OpenStudy (ashsmashum1290):

What is the area of the circle? use 3.14. A.12.56 mm B.25.12 mm C.50.24 mm D.502.4 mm

OpenStudy (ashsmashum1290):



jagr2713 (jagr2713):

Ok the formula for area of circle is \[A=\pi \ r ^{2}\] when you plug it in what do you get @Ashsmashum1290

OpenStudy (ashsmashum1290):

12.56?

OpenStudy (ashsmashum1290):

4

jagr2713 (jagr2713):

Good now what did you do wrong with 12.56 answer?

OpenStudy (ashsmashum1290):

i did 3.14x4

jagr2713 (jagr2713):

you forgetting something @Ashsmashum1290 look at the formula again \[A=\pi \ r ^{2}\]

jagr2713 (jagr2713):

dont forget the r^2 thats a common mistake made by people but costly. Thats why they put a distractor in the answer choices that looks like the answer but is not.

OpenStudy (ashsmashum1290):

what do i do divide?

jagr2713 (jagr2713):

No remeber when you square something you *Multiply* it so if its 12^2 that means its 12 times 12 which is 144 or 5^2=25 you get it? @Ashsmashum1290

jagr2713 (jagr2713):

so if 4^2 what is it?

OpenStudy (ashsmashum1290):

8

jagr2713 (jagr2713):

Thats 4*2 i mean 4 square remember what i said above. If it was 12 times 2 it would of been 24 but its 12 square which is 144. In other words its like 4*4 or 12*12 or 15*15

jagr2713 (jagr2713):

theres square and cubed. Cubed is by 3. so if its 12^3 or cubed it is 1728

jagr2713 (jagr2713):

Lets try one more time. You got this :D 4^2=? @Ashsmashum1290

OpenStudy (ashsmashum1290):

what is that sign in the middle of the 4 and 2?????

jagr2713 (jagr2713):

\[4^{2}=?\]

OpenStudy (ashsmashum1290):

4x2=8

jagr2713 (jagr2713):

no D: read what i said above..

jagr2713 (jagr2713):

Thats 4*2 i mean 4 square remember what i said above. If it was 12 times 2 it would of been 24 but its 12 square which is 144. In other words its like 4*4 or 12*12 or 15*15

jagr2713 (jagr2713):

Please read it

OpenStudy (ashsmashum1290):

okay i read it but i still don't get it :(

jagr2713 (jagr2713):

ok you know how 4*2=8 right? so for this it squared which is like saying the number multiplied by its self so for example 5*2=10 but \[5^{2}=25\]

OpenStudy (ashsmashum1290):

16?

jagr2713 (jagr2713):

Correct so if you plug it in what do you get \[A=\pi \ 4^{2}\]

OpenStudy (ashsmashum1290):

plug in the 3.14?

jagr2713 (jagr2713):

yea :D

OpenStudy (ashsmashum1290):

50.24

jagr2713 (jagr2713):

the \[\pi \ is \ the \ 3.14\]

jagr2713 (jagr2713):

Correct :D

jagr2713 (jagr2713):

See was that hard @Ashsmashum1290

OpenStudy (ashsmashum1290):

okay thanks a lot i never would have figured that out :)

jagr2713 (jagr2713):

Anytime LD

jagr2713 (jagr2713):

:D